The resistance of a uniform copper wire 50.0 meters long and 1.15 mm in

diameter is 0.830 ohms at

20° C. What is the resistivity of the copper at this temperature?


1
Expert's answer
2022-05-20T13:19:35-0400
"R=\rho\frac{l}{A}"

"\rho=RA/l=R\pi(d/2)^2/l\\
=0.830*3.14*(0.0015/2)^2/50.0=2.93*10^{-8}\:\rm \Omega\cdot m"